# Communication-Efficient Stochastic Distributed Learning

### v1

@@ A-C Deviation from the average
$\begin{bmatrix}\widehat{\mathbf{Q}}^{T}\mathbf{X}_{k}\\ \widehat{\mathbf{Q}}^{T}\mathbf{Y}_{k}\\ \widehat{\mathbf{Q}}^{T}\tilde{\mathbf{Y}}_{k}.\end{bmatrix}=\boldsymbol{\phi}^{T}\left[\begin{array}[]{ccc}-\gamma\tau\mathbf{I}&d_{12}\mathbf{I}&d_{13}\mathbf{I}\\ \mathbf{I}&d_{22}\mathbf{I}&d_{23}\mathbf{I}\\ \mathbf{I}&\mathbf{I}&\mathbf{I}\end{array}\right]\mathbf{P}_{0}^{T}\widehat{\mathbf{d}}_{k}$
$=\boldsymbol{\phi}^{T}\left[\begin{array}[]{c}-\gamma\tau\mathbf{I}\mathbf{P}_{0}^{T}[1]+d_{12}\mathbf{P}_{0}^{T}[2]+d_{13}\mathbf{P}_{0}^{T}[3]\\ \mathbf{P}_{0}^{T}[1]+d_{22}\mathbf{P}_{0}^{T}[2]+d_{23}\mathbf{P}_{0}^{T}[3]\\ \mathbf{P}_{0}^{T}[1]+\mathbf{P}_{0}^{T}[2]+\mathbf{P}_{0}^{T}[3]\end{array}\right]\widehat{\mathbf{d}}_{k},$
where $\mathbf{P}_{0}^{T}[1],\mathbf{P}_{0}^{T}[2],\mathbf{P}_{0}^{T}[3]$ are the top, middle and bottom blocks of $\mathbf{P}_{0}^{T}$ respectively. Moreover, we have $|d_{12}|^{2}=|d_{13}|^{2}\leq\frac{2\gamma\tau}{|\tilde{\lambda}_{\max}|\rho}$, $|d_{22}|=|d_{23}|=1$. Now, if we let $\gamma\tau\leq\frac{2}{|\tilde{\lambda}_{\max}|\rho},$ and using $\|\phi\|=1$, $\|\mathbf{P}_{0}^{T}[i]\|=1$, $i=1,2,3$, we derive that
@@
$\quad=\|\phi^{T}(-\gamma\tau\mathbf{I}\mathbf{P}_{0}^{T}[1]+d_{12}\mathbf{P}_{0}^{T}[2]+d_{13}\mathbf{P}_{0}^{T}[3])\widehat{\mathbf{d}}_{k}\|^{2}$
$\quad\leq 3(\gamma^{2}\tau^{2}+\frac{4\gamma\tau}{|\tilde{\lambda}_{\max}|\rho})\|\widehat{\mathbf{d}}_{k}\|^{2}\leq\frac{18\gamma\tau}{|\tilde{\lambda}_{\max}|\rho}\|\widehat{\mathbf{d}}_{k}\|^{2}.$
@@
$\|\sum_{t=0}^{\tau-1}(G({\Phi}_{k}^{t})-\nabla\mathrm{F}(\bar{\mathbf{X}}_{k})+\rho\mathbf{D}{\Phi}_{k}^{t}-\rho\mathbf{D}\mathbf{X}_{k})\|^{2}$
$=\|\sum_{t=0}^{\tau-1}(G({\Phi}_{k}^{t})-\nabla F(\mathbf{\Phi}_{k}^{t})+\nabla F({\Phi}_{k}^{t})-\nabla\mathrm{F}(\bar{\mathbf{X}}_{k})$
$+\rho\mathbf{D}({\Phi}_{k}^{t}-\bar{\mathbf{X}}_{k}+\bar{\mathbf{X}}_{k}-\mathbf{X}_{k}))\|^{2}$
$\leq 4(\tau L^{2}+\tau\rho^{2}d_{u}^{2})\|\widehat{\mathbf{\Phi}}_{k}\|^{2}+4\tau^{2}\rho^{2}d_{u}^{2}\left\|\bar{\mathbf{X}}_{k}-\mathbf{X}_{k}\right\|^{2}$
$+4\tau\sum_{t=0}^{\tau-1}\|G({\Phi}_{k}^{t})-\nabla F({\Phi}_{k}^{t})\|^{2}.$
Denoting $\epsilon_{k}=\sum_{t=0}^{\tau-1}(\rho\mathbf{1}^{T}\mathbf{D}\Phi_{k}^{t}-\mathbf{1}^{T}\mathbf{A}^{T}\mathbf{Z}_{k})=\rho\sum_{t=0}^{\tau-1}\sum_{i=1}^{N}(d_{i}\phi_{i,k}^{t}-d_{i}\bar{x}_{k}+d_{i}\bar{x}_{k}-d_{i}x_{i,k})$, we have
$\epsilon_{k}^{2}\leq 2\rho^{2}d_{u}^{2}\tau N(\|\widehat{\mathbf{\Phi}}_{k}\|^{2}+\tau\|\mathbf{X}_{k}-\bar{\mathbf{X}}_{k}\|^{2}).$ (27)
Now, denoting $\overline{G}(\boldsymbol{\Phi}^{t}_{k})=\frac{1}{N}\sum_{i}g_{i}(\phi_{i,k}^{t})$, then $\|\sum_{t}\overline{G}(\boldsymbol{\Phi}^{t}_{k})\|^{2}\leq 2\|\sum_{t}\overline{\nabla F}(\boldsymbol{\Phi}^{t}_{k})\|^{2}+\frac{2\tau}{N}\sum_{t}\sum_{i}\|\nabla f_{i}(\phi_{i,k}^{t})-g_{i}(\phi_{i,k}^{t})\|^{2},$ where $\|\sum_{t}\overline{\nabla F}(\boldsymbol{\Phi}^{t}_{k})\|^{2}=\|\frac{1}{N}\sum_{i}\sum_{t}\nabla f_{i}\left(\phi_{i,k}^{t}\right)\|^{2}\leq\frac{2\tau}{N}\sum_{i}\sum_{t}\|\nabla f_{i}(\phi_{i,k}^{t})-\nabla f_{i}(\bar{x}_{k})\|^{2}+2\tau^{2}\|\frac{1}{N}\sum_{i}\nabla f_{i}(\bar{x}_{k})\|^{2}.$ We also have
$\left\|\nabla F(\bar{\mathbf{X}}_{k+1})-\nabla F(\bar{\mathbf{X}}_{k})\right\|^{2}=N\left\|\nabla F(\bar{x}_{k+1})-\nabla F(\bar{x}_{k})\right\|^{2}$
$\leq NL^{2}\|\frac{\gamma}{N}\sum_{t=0}^{\tau-1}\sum_{i=1}^{N}g_{i}(\phi_{k}^{t,i})+\frac{\gamma}{N}\sum_{t=0}^{\tau-1}(\rho\mathbf{1}^{T}\mathbf{D}\Phi_{k}^{t}-\mathbf{1}^{T}\mathbf{A}^{T}\mathbf{Z}_{k})\|^{2}$
$\leq 2L^{2}\gamma^{2}\left(2\tau\sum_{t}\sum_{i}\|\nabla f_{i}\left(\phi_{i,k}^{t}\right)-g_{i}\left(\phi_{i,k}^{t}\right)\|^{2}\right.$
$\left.+2N\|\sum_{t}\overline{\nabla F}\left(\boldsymbol{\Phi}^{t}_{k}\right)\|^{2}+2\rho^{2}d_{u}^{2}\tau(\|\widehat{\mathbf{\Phi}}_{k}\|^{2}+\tau\|\mathbf{X}_{k}-\bar{\mathbf{X}}_{k}\|^{2})\right).$
Using $\|\widehat{\mathbf{Q}}^{T}\|=1$ it further holds that:
$\|\mathbf{\widehat{h}}_{k}\|^{2}\leq\gamma^{2}(1+2\rho^{2}\|\tilde{\mathbf{L}}\|^{2})\|\widehat{\mathbf{V}}^{-1}\|^{2}\cdot$
$\cdot\|\sum_{t=0}^{\tau-1}(G(\Phi_{k}^{t})-\nabla\mathrm{F}(\bar{\mathbf{X}}_{k})+\rho\mathbf{D}\Phi_{k}^{t}-\rho\mathbf{D}\mathbf{X}_{k})\|^{2}$
$+3\|\widehat{\mathbf{V}}^{-1}\|^{2}\left\|\nabla F(\bar{\mathbf{X}}_{k+1})-\nabla F(\bar{\mathbf{X}}_{k})\right\|^{2}$
$\leq\gamma^{2}(1+2\rho^{2}\|\tilde{\mathbf{L}}\|^{2})\|\widehat{\mathbf{V}}^{-1}\|^{2}\left(4(\tau L^{2}+\tau\rho^{2}d_{u}^{2})\|\widehat{\mathbf{\Phi}}_{k}\|^{2}\right.$
$\left.+4\tau^{2}\rho^{2}d_{u}^{2}\left\|\bar{\mathbf{X}}_{k}-\mathbf{X}_{k}\right\|^{2}+4\tau\sum_{t=0}^{\tau-1}\|G(\Phi_{k}^{t})-\nabla F(\Phi_{k}^{t})\|^{2}\right)$
$+{6L^{2}\gamma^{2}}\|\widehat{\mathbf{V}}^{-1}\|^{2}\cdot\left(2\tau\sum_{t}\sum_{i}\|\nabla f_{i}\left(\phi_{i,k}^{t}\right)-g_{i}\left(\phi_{i,k}^{t}\right)\|^{2}\right.$
$\left.+2N\|\sum_{t}\overline{\nabla F}\left({\Phi}^{t}_{k}\right)\|^{2}+2\rho^{2}d_{u}^{2}\tau(\|\widehat{\mathbf{\Phi}}_{k}\|^{2}+\tau\|\mathbf{X}_{k}-\bar{\mathbf{X}}_{k}\|^{2})\right).$ (28)
@@
Recalling (26), using Jensen’s inequality $\|\widehat{\mathbf{d}}_{k+1}\|^{2}\leq\frac{1}{\|\mathbf{\Delta}\|}\|\mathbf{\Delta}\|^{2}\|\widehat{\mathbf{d}}_{k}\|^{2}+\frac{1}{1-\|\mathbf{\Delta}\|}\|\mathbf{\widehat{h}}_{k}\|^{2}$ yields (17). ∎
Appendix B Convergence analysis for LT-ADMM
B-A Key bounds
Lemma 2.
@@ B-A Key bounds
Let Assumptions 1, 2, and 4 hold; when $\gamma<\frac{2}{\tau|\tilde{\lambda}_{\max}|\rho}$ and $4\gamma^{2}\tau(L^{2}+\rho^{2}d_{u}^{2})<\frac{1/4}{\tau-1}$, we have
$\mathbb{E}\left[\|\widehat{\mathbf{\Phi}}_{k}\|^{2}\right]$ (29)
@@
$\leq\left(\frac{72\gamma\tau^{2}}{|\tilde{\lambda}_{\max}|\rho}+(18+\frac{36\gamma\tau\rho^{2}d_{u}^{2}}{|\tilde{\lambda}_{\max}|\rho})16\tau^{3}\gamma^{2}\right)\mathbb{E}\left[\|\widehat{\mathbf{d}}_{k}\|^{2}\right]$
$+16\tau^{3}\gamma^{2}N\mathbb{E}\left\|\nabla F(\bar{x}_{k})\right\|^{2}+4N\tau^{2}\gamma^{2}\sigma^{2}.$
Proof.
@@
$-\frac{\gamma}{N}\left(\sum_{t=0}^{\tau-1}(\rho\mathbf{1}^{T}\mathbf{D}\Phi_{k}^{t}-\mathbf{1}^{T}\mathbf{A}^{T}\mathbf{Z}_{k})\right)$
@@
$\Phi_{k}^{t+1}=\Phi_{k}^{t}+\gamma\mathbf{Y}_{k}-\gamma(G(\Phi_{k}^{t})-\nabla\mathrm{F}(\bar{\mathbf{X}}_{k})+\rho\mathbf{D}\Phi_{k}^{t}-\rho\mathbf{D}\mathbf{X}_{k})$ (31)
Recall that by Assumption 4, $\|G(\Phi_{k}^{t})-\nabla F(\Phi_{k}^{t})\|^{2}\leq N\sigma^{2}$.
Now, suppose that $\tau\geq 2$, using Jensen’s inequality we obtain
@@
$\leq\mathbb{E}[\|\Phi_{k}^{t}-\bar{\mathbf{X}}_{k}+\gamma\mathbf{Y}_{k}-\gamma(\nabla F(\Phi_{k}^{t})-\nabla\mathrm{F}(\bar{\mathbf{X}}_{k})+$
$+\rho\mathbf{D}\Phi_{k}^{t}-\rho\mathbf{D}\mathbf{X}_{k})\|^{2}]+N\gamma^{2}\sigma^{2}$
$\leq(1+\frac{1}{\tau-1})\mathbb{E}[\|\Phi_{k}^{t}-\bar{\mathbf{X}}_{k}\|^{2}]+N\gamma^{2}\sigma^{2}+$
$+\tau\gamma^{2}\mathbb{E}[\|\mathbf{Y}_{k}-(\nabla F(\Phi_{k}^{t})-\nabla\mathrm{F}(\bar{\mathbf{X}}_{k})+\rho\mathbf{D}\Phi_{k}^{t}-\rho\mathbf{D}\mathbf{X}_{k})\|^{2}]$
${\leq}(1+\frac{1}{\tau-1})\mathbb{E}[\|\Phi_{k}^{t}-\bar{\mathbf{X}}_{k}\|^{2}]+N\gamma^{2}\sigma^{2}+$
$+2\tau\gamma^{2}\mathbb{E}[\|\nabla F(\Phi_{k}^{t})-\nabla\mathrm{F}(\bar{\mathbf{X}}_{k})+\rho\mathbf{D}\Phi_{k}^{t}-\rho\mathbf{D}\bar{\mathbf{X}}_{k}\|^{2}]+$
$+2\tau\gamma^{2}\mathbb{E}[\left\|\mathbf{Y}_{k}-\bar{\mathbf{Y}}_{k}+\bar{\mathbf{Y}}_{k}+(\rho\mathbf{D}\bar{\mathbf{X}}_{k}-\rho\mathbf{D}\mathbf{X}_{k})\right\|^{2}]$
$\leq(1+\frac{1}{\tau-1}+4\gamma^{2}\tau(L^{2}+\rho^{2}d_{u}^{2}))\mathbb{E}[\|\Phi_{k}^{t}-\bar{\mathbf{X}}_{k}\|^{2}+$
$+2\tau\gamma^{2}\left\|\mathbf{Y}_{k}-\bar{\mathbf{Y}}_{k}+\bar{\mathbf{Y}}_{k}+(\rho\mathbf{D}\bar{\mathbf{X}}_{k}-\rho\mathbf{D}\mathbf{X}_{k})\right\|^{2}]+N\gamma^{2}\sigma^{2}$
$\leq(1+\frac{5/4}{\tau-1})\mathbb{E}[\left\|\Phi_{k}^{t}-\bar{\mathbf{X}}_{k}\right\|^{2}]+N\gamma^{2}\sigma^{2}+$
$+2\tau\gamma^{2}\mathbb{E}[\|\mathbf{Y}_{k}-\bar{\mathbf{Y}}_{k}+\bar{\mathbf{Y}}_{k}+(\rho\mathbf{D}\bar{\mathbf{X}}_{k}-\rho\mathbf{D}\mathbf{X}_{k})\|^{2}],$
where $\bar{\mathbf{Y}}_{k}=\mathbf{1}\otimes\frac{1}{N}\mathbf{1}^{T}\nabla\mathrm{F}(\bar{\mathbf{X}}_{k})=\mathbf{1}\otimes\frac{1}{N}\sum_{i}\nabla f_{i}(\bar{x}_{k})$, and the last inequality holds when
$4\gamma^{2}\tau(L^{2}+\rho^{2}d_{u}^{2})<\frac{1/4}{\tau-1}.$ (32)
Iterating the above inequality for $t=0,...,\tau-1$
@@
$+2\tau\gamma^{2}\sum_{l=0}^{t}\left(1+\frac{5/4}{\tau-1}\right)^{l}\cdot$
$\cdot\mathbb{E}[\left\|\mathbf{Y}_{k}-\bar{\mathbf{Y}}_{k}+\bar{\mathbf{Y}}_{k}+(\rho\mathbf{D}\bar{\mathbf{X}}_{k}-\rho\mathbf{D}\mathbf{X}_{k})\right\|^{2}]+$
@@
$\leq 4\mathbb{E}[\left\|\mathbf{X}_{k}-\bar{\mathbf{X}}_{k}\right\|^{2}]+4\tau N\gamma^{2}\sigma^{2}+$
$+8\tau^{2}\gamma^{2}\mathbb{E}[\left\|\mathbf{Y}_{k}-\bar{\mathbf{Y}}_{k}+\bar{\mathbf{Y}}_{k}+(\rho\mathbf{D}\bar{\mathbf{X}}_{k}-\rho\mathbf{D}\mathbf{X}_{k})\right\|^{2}],$
where the last inequality holds by $(1+\frac{a}{\tau-1})^{t}\leq\exp(\frac{at}{\tau-1})\leq\exp(a)$ for $t\leq\tau-1$ and $a={5}/{4}$.
@@
$\begin{split}&\mathbb{E}[\|\widehat{\mathbf{\Phi}}_{k}\|^{2}]\leq 4\tau\mathbb{E}[\|\mathbf{X}_{k}-\bar{\mathbf{X}}_{k}\|^{2}]+16\tau^{3}\gamma^{2}N\mathbb{E}[\left\|\nabla F(\bar{x}_{k})\right\|^{2}]+\\ &+4N\tau^{2}\gamma^{2}\sigma^{2}+16\tau^{3}\gamma^{2}\mathbb{E}[\|\mathbf{Y}_{k}-\bar{\mathbf{Y}}_{k}+(\rho\mathbf{D}\bar{\mathbf{X}}_{k}-\rho\mathbf{D}\mathbf{X}_{k})\|^{2}];\end{split}$ (33)
moreover, it is easy to verify that (33) also holds for $\tau=1$. As a consequence of (16), we obtain
$\left\|\mathbf{Y}_{k}-\bar{\mathbf{Y}}_{k}+(\rho\mathbf{D}\bar{\mathbf{X}}_{k}-\rho\mathbf{D}\mathbf{X}_{k})\right\|^{2}\leq(18+\frac{36\gamma\tau\rho^{2}d_{u}^{2}}{|\tilde{\lambda}_{\max}|\rho})\|\widehat{\mathbf{d}}_{k}\|^{2},$ (34)
and using this fact into (33) concludes the proof. ∎
Lemma 3.
Let Assumptions 1, 2, and 4 hold; when $\gamma\leq\min\{\bar{\gamma}_{1},\bar{\gamma}_{2}\}$, it holds that
$\mathbb{E}[\|\widehat{\mathbf{d}}_{k+1}\|^{2}]\leq(\delta+\frac{q_{1}}{1-\delta})\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]+$ (35)
$+\frac{q_{2}}{1-\delta}\mathbb{E}[\|\sum_{t}\overline{\nabla F}(\boldsymbol{\Phi}^{t}_{k})\|^{2}]+\frac{q_{3}}{1-\delta}\mathbb{E}[\|\nabla F(\bar{x}_{k})\|^{2}]+q_{4}\sigma^{2}$

### v2

@@ 7.3 Deviation from the average
$\begin{bmatrix}\widehat{\mathbf{Q}}^{T}\mathbf{X}_{k}\\ \widehat{\mathbf{Q}}^{T}\mathbf{Y}_{k}\\ \widehat{\mathbf{Q}}^{T}\tilde{\mathbf{Y}}_{k}.\end{bmatrix}=\boldsymbol{\phi}^{T}\left[\begin{array}[]{ccc}-\beta\tau\mathbf{I}&d_{12}\mathbf{I}&d_{13}\mathbf{I}\\ \mathbf{I}&d_{22}\mathbf{I}&d_{23}\mathbf{I}\\ \mathbf{I}&\mathbf{I}&\mathbf{I}\end{array}\right]\mathbf{P}_{0}^{T}\widehat{\mathbf{d}}_{k}$
$=\boldsymbol{\phi}^{T}\left[\begin{array}[]{c}-\beta\tau\mathbf{I}\mathbf{P}_{0}^{T}[1]+d_{12}\mathbf{P}_{0}^{T}[2]+d_{13}\mathbf{P}_{0}^{T}[3]\\ \mathbf{P}_{0}^{T}[1]+d_{22}\mathbf{P}_{0}^{T}[2]+d_{23}\mathbf{P}_{0}^{T}[3]\\ \mathbf{P}_{0}^{T}[1]+\mathbf{P}_{0}^{T}[2]+\mathbf{P}_{0}^{T}[3]\end{array}\right]\widehat{\mathbf{d}}_{k},$
where $\mathbf{P}_{0}^{T}[1],\mathbf{P}_{0}^{T}[2],\mathbf{P}_{0}^{T}[3]$ are the top, middle and bottom blocks of $\mathbf{P}_{0}^{T}$ respectively. Moreover, we have $|d_{12}|^{2}=|d_{13}|^{2}\leq\frac{2\beta\tau}{\lambda_{l}\rho}$, $|d_{22}|=|d_{23}|=1$. Now, if we let $\beta\tau\leq\frac{2}{\lambda_{l}\rho},$ and using $\|\phi\|=1$, $\|\mathbf{P}_{0}^{T}[i]\|=1$, $i=1,2,3$, we derive that
@@
$\quad=\|\phi^{T}(-\beta\tau\mathbf{I}\mathbf{P}_{0}^{T}[1]+d_{12}\mathbf{P}_{0}^{T}[2]+d_{13}\mathbf{P}_{0}^{T}[3])\widehat{\mathbf{d}}_{k}\|^{2}$
$\quad\leq 3(\beta^{2}\tau^{2}+\frac{4\beta\tau}{\lambda_{l}\rho})\|\widehat{\mathbf{d}}_{k}\|^{2}\leq\frac{18\beta\tau}{\lambda_{l}\rho}\|\widehat{\mathbf{d}}_{k}\|^{2}.$
@@
$\|\sum_{t=0}^{\tau-1}(G(\mathrm{\Phi}_{k}^{t})-\nabla\mathrm{F}(\bar{\mathbf{X}}_{k}))\|^{2}$
$\leq 2\tau L^{2}\|\widehat{\mathbf{\Phi}}_{k}\|^{2}+2\tau\sum_{t=0}^{\tau-1}\|G(\mathrm{\Phi}_{k}^{t})-\nabla F(\mathrm{\Phi}_{k}^{t})\|^{2}.$
Denote $\overline{G}(\mathrm{\Phi}_{k}^{t})=\frac{1}{N}\sum_{i=1}^{N}g_{i}(\phi_{i,k}^{t})$ and $\overline{\nabla\mathrm{F}}(\mathrm{\Phi}_{k}^{t})=\frac{1}{N}\sum_{i=1}^{N}\nabla f_{i}(\phi_{i,k}^{t})$, we have
$\|\sum_{t=0}^{\tau-1}\overline{G}(\mathrm{\Phi}_{k}^{t})\|^{2}$ (27)
$=\|\frac{1}{N}\sum_{i}\sum_{t}(\nabla f_{i}\left(\phi_{i,k}^{t}\right)+g_{i}\left(\phi_{i,k}^{t}\right)-\nabla f_{i}\left(\phi_{i,k}^{t}\right))\|^{2}$
$\leq 2\|\sum_{t=0}^{\tau-1}\overline{\nabla\mathrm{F}}(\mathrm{\Phi}_{k}^{t})\|^{2}+\frac{2\tau}{N}\sum_{i}\sum_{t}\|g_{i}\left(\phi_{i,k}^{t}\right)-\nabla f_{i}\left(\phi_{i,k}^{t}\right)\|^{2}$
We also have $\|\nabla F(\bar{\mathbf{X}}_{k+1})-\nabla F(\bar{\mathbf{X}}_{k})\|^{2}=NL^{2}\left\|\bar{x}_{k+1}-\bar{x}_{k}\right\|^{2}=NL^{2}\gamma^{2}\|\sum_{t}\overline{G}(\mathrm{\Phi}_{k}^{t})\|^{2}$, it further holds that:
$\|{\mathbf{h}}_{k}\|^{2}\leq\gamma^{2}(1+2\rho^{2}\|\tilde{\mathbf{L}}\|^{2})(2\tau L^{2}\|\widehat{\mathbf{\Phi}}_{k}\|^{2}$ (28)
$+2\tau\sum_{t=0}^{\tau-1}\|G(\mathrm{\Phi}_{k}^{t})-\nabla F(\mathrm{\Phi}_{k}^{t})\|^{2})$
$+6L^{2}\frac{\gamma^{4}}{\beta^{2}}\bigg(\tau\sum_{i}\sum_{t}\left\|g_{i}\left(\phi_{i,k}^{t}\right)-\nabla f_{i}\left(\phi_{i,k}^{t}\right)\right\|^{2}$
$+N\|\sum_{t=0}^{\tau-1}\overline{\nabla\mathrm{F}}(\mathrm{\Phi}_{k}^{t})\|^{2}\bigg)$
@@
Recalling (26), using Jensen’s inequality $\|\widehat{\mathbf{d}}_{k+1}\|^{2}\leq\frac{1}{\|\mathbf{\Delta}\|}\|\mathbf{\Delta}\|^{2}\|\widehat{\mathbf{d}}_{k}\|^{2}+\frac{1}{1-\|\mathbf{\Delta}\|}\|\mathbf{\widehat{h}}_{k}\|^{2}$ yields (17).
8 Convergence analysis for LT-ADMM
8.1 Key bounds
Lemma 8.4.
@@ 8.1 Key bounds
Let Assumptions 1, 2, and 4 hold; when $\beta<\frac{2}{\tau\lambda_{u}\rho}$ and $\gamma\leq\bar{\gamma}_{1}$, we have
@@
$\mathbb{E}\left[\|\widehat{\mathbf{\Phi}}_{k}\|^{2}\right]$ $\leq\left(\frac{72\beta\tau^{2}}{\lambda_{l}\rho}+144\tau^{3}\beta^{2}\right)\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]+4N\tau^{2}\gamma^{2}\sigma^{2}$ (29)
[…]
Proof 8.5.
@@
$\Phi_{k}^{t+1}=\Phi_{k}^{t}+\beta\mathbf{Y}_{k}-\gamma(G(\Phi_{k}^{t})-\nabla\mathrm{F}(\bar{\mathbf{X}}_{k}))$ (31)
Recall that by Assumption 4, $\|G(\Phi_{k}^{t})-\nabla F(\Phi_{k}^{t})\|^{2}\leq N\sigma^{2}$. Now, suppose that $\tau\geq 2$, using Jensen’s inequality we obtain
@@
$=\mathbb{E}[\|\Phi_{k}^{t}-\bar{\mathbf{X}}_{k}+\beta\mathbf{Y}_{k}-\gamma(\nabla\mathrm{F}(\mathrm{\Phi}_{k}^{t})-\nabla\mathrm{F}(\bar{\mathbf{X}}_{k}))\|^{2}]+N\gamma^{2}\sigma^{2}$
$\leq\left(1+\frac{1}{\tau-1}\right)\mathbb{E}[\left\|\Phi_{k}^{t}-\bar{\mathbf{X}}_{k}\right\|^{2}]+N\gamma^{2}\sigma^{2}$ (32)
$+\tau\mathbb{E}[\|\beta Y_{k}-\gamma(\nabla\mathrm{F}(\mathrm{\Phi}_{k}^{t})-\nabla\mathrm{F}(\bar{\mathbf{X}}_{k}))\|^{2}]$
$\leq\left(1+\frac{1}{\tau-1}+2\gamma^{2}\tau L^{2}\right)\mathbb{E}[\left\|\mathbf{\Phi}_{k}^{t}-\bar{\mathbf{X}}_{k}\right\|^{2}]$ (33)
$+2\tau\beta^{2}\mathbb{E}[\left\|\mathbf{Y}_{k}\right\|^{2}]+\gamma^{2}N\sigma^{2}$
$\leq\left(1+\frac{5/4}{\tau-1}\right)\mathbb{E}[\|\mathbf{\Phi}_{k}^{t}-\bar{\mathbf{X}}_{k}\|^{2}]+\gamma^{2}N\sigma^{2}+2\tau\beta^{2}\mathbb{E}[\|\mathbf{Y}_{k}\|^{2}],$ (34)
where the last inequality holds when
$2\gamma^{2}\tau L^{2}\leq\frac{1/4}{\tau-1},$ (35)
which is satisfied by $\gamma\leq\bar{\gamma}_{1}$. Iterating the above inequality for $t=0,...,\tau-1$
@@
$+2\tau\beta^{2}\sum_{l=0}^{t}\left(1+\frac{5/4}{\tau-1}\right)^{l}\mathbb{E}[\left\|\mathbf{Y}_{k}-\bar{\mathbf{Y}}_{k}+\bar{\mathbf{Y}}_{k}\right\|^{2}]$
@@
$\leq 4\mathbb{E}[\left\|\mathbf{X}_{k}-\bar{\mathbf{X}}_{k}\right\|^{2}]+4\tau N\gamma^{2}\sigma^{2}+8\tau^{2}\beta^{2}\mathbb{E}[\left\|\mathbf{Y}_{k}-\bar{\mathbf{Y}}_{k}+\bar{\mathbf{Y}}_{k}\right\|^{2}],$
where the last inequality holds by $(1+\frac{a}{\tau-1})^{t}\leq\exp(\frac{at}{\tau-1})\leq\exp(a)$ for $t\leq\tau-1$ and a=5/4a={5}/{4}.
@@
$\begin{split}\mathbb{E}[\|\widehat{\mathbf{\Phi}}_{k}\|^{2}]&\leq 4\tau\mathbb{E}[\|\mathbf{X}_{k}-\bar{\mathbf{X}}_{k}\|^{2}]+4N\tau^{2}\gamma^{2}\sigma^{2}\\ &+16\tau^{3}\beta^{2}\mathbb{E}[\|\mathbf{Y}_{k}-\bar{\mathbf{Y}}_{k}\|^{2}]+16\tau^{3}N\gamma^{2}\mathbb{E}[\|\nabla F(\bar{x}_{k})\|^{2}];\end{split}$ (36)
moreover, it is easy to verify that (36) also holds for $\tau=1$. Using (16) concludes the proof.
Lemma 8.6.
Let Assumptions 1, 2, and 4 hold. When $\beta<\frac{2}{\tau\lambda_{u}\rho}$ and $\gamma\leq\bar{\gamma}_{1}$,
$\mathbb{E}[\|\widehat{\mathbf{d}}_{k+1}\|^{2}]$ (37)
$\leq(\delta+\frac{c_{0}}{1-\delta})\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]+\frac{c_{1}}{1-\delta}\mathbb{E}[\|\sum_{t}\overline{\nabla F}(\boldsymbol{\Phi}^{t}_{k})\|^{2}]$
$+\frac{c_{2}}{1-\delta}\mathbb{E}[\|\nabla F(\bar{x}_{k})\|^{2}]+\frac{c_{3}}{1-\delta}\sigma^{2}$